HUBBLE SPACE TELESCOPE DAILY REPORT #2822 PERIOD COVERED: 0000Z (UTC) 03/05/01 - 0000Z (UTC) 03/06/01 Daily Status Report as of 065/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of STIS/CCD 8562 (Probing the Large Scale Structure: Cosmic Shear Observations) The Space Telescope Imaging Spectrograph (CCD) was used to probe the distortion of light bundles from distant galaxies, looking at the statistical properties of the intervening inhomogeneous {dark} matter distribution. The proposal completed nominally. 1.2 Completed WF/PC-2 8829 (Observatory Focus Monitor) The WF/PC-2 was used to analyze the HST focus that drifts slowly and shows evidence of undergoing slips of a few microns at random times. The rate of the WFPC2 monitoring program is insufficient to track and/or understand OTA behavior in order to request timely and appropriate secondary mirror corrections. This 14-orbit program obtains a large amount of high signal to noise focus data, sometimes in two Science Instruments at once, and will be sufficient to more accurately define the HST focus. There were no reported problems. 1.3 Completed WF/PC-2 8175 (A Survey of Gravitational Lenses as Cosmological Tools III) The WF/PC-2 was used to make observations of the gravitational lens system Q0252-3249. Gravitational lenses are powerful tools for studying cosmology, galactic structure, and galaxy evolution. The observations were completed as planned, and no anomalies were reported. 1.4 Completed STIS/CCD 8663 (Survey of SMC Planetary Nebulae: Nebular and Stellar Evolution in a Low- Metallicity Environment) The Space Telescope Imaging Spectrograph (CCD) was used to survey of SMC planetary nebulae {PNe} in order to study the co-evolution of the nebulae and their central stars, in an environment that is chemically very metal poor. We will obtain STIS imaging and medium-resolution slitless spectroscopy that will yield line fluxes and nebular morphologies in important emission lines, plus magnitudes of the central stars. From these data we will gather a harvest of information: the nebular size, morphology, ionization structure, density, and mass; and the central star temperature, luminosity, and mass. The proposal completed with no reported anomalies. 1.5 Completed WF/PC-2 8805 (POMS Test Proposal: WFII Parallel Archive Proposal Continuation) The WF/PC-2 was used to perform a generic target version of the WFPC2 Archival Pure Parallel program. The program will be used to take parallel images of random areas of the sky, following the recommendations of the Parallels Working Group. The proposal completed with no reported problems. 1.6 Completed WF/PC-2 8827 (Cycle 9 Supplemental Darks pt2/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The proposal completed nominally. 1.7 Completed WF/PC-2 8583 (Imaging Snapshots of Asteroids) The WF/PC-2 was used to obtain images of the fifty largest main belt asteroids that have favorable apparitions during cycle 9. The images will be searched for companion bodies, as well as mineralogical variegation on the resolved main bodies. There were no reported problems. * * * * * * * * * * * * * * * * * * * * * * * * * * * * * * * * * * * * The following proposals, scheduled to be executed during this operational day, were not executed due to the entry into safemode. * * * * * * * * * * * * * * * * * * * * * * * * * * * * * * * * * * * * Three Sets of STIS/CCD 8562 (Probing the Large Scale Structure: Cosmic Shear Observations) WF/PC-2 8805 (POMS Test Proposal: WFII Parallel Archive Proposal Continuation) Two Sets of WF/PC-2 8827 (Cycle 9 Supplemental Darks pt2/3) WF/PC-2 8583 (Imaging Snapshots of Asteroids) Two Sets of WF/PC-2 8723 (A Snapshot Survey of Galactic Globular Clusters) Two Sets of STIS/CCD 8864 (CCD Dark Monitor-Part 2) WF/PC-2 8632 (A UV Atlas of Nearby Galaxies) STIS/CCD 8687 (Elliptical Galaxies With Nuclear Disks of Stars: Black Hole Search and Stellar Populations) Five Sets of WF/PC-2 8811 (Cycle 9 Standard Darks) WF/PC-2 8812 (Cycle 9 Internal Monitor) STIS/CCD 8865 (Bias Monitor-Part 2) STIS/CCD/MA1/MA1 8623 (HST, Chandra, and FUSE Studies of Interstellar Material toward HD 24534 {aka X Persei}) WF/PC-2 8817 (Cycle 9 INTFLAT Sweeps and Linearity Test) STIS/CCD/MA1/MA2 8857 (MAMA Sensitivity and Focus Monitor C9) 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 0 Successful: 0 Scheduled Re-acquisitions: 0 Successful: 0 2.2 FHST Updates: Scheduled: 0 Successful: 0 2.3 Operations Notes: The HST entered Software Sunpoint (SWSP) at 064/063856Z due to the failure of the Magnetic Torquer Electronics #1 (MCE-1) command echo check. The following actions, decided upon and coordinated at meetings in the STOCC area and executed in accordance with operations requests, were performed in the recovery process: The GEA was powered up at 064/1420Z. The programmable telemetry format was synchronized to the "top of the minute" at 064/1530Z. Engineering status buffer and selected dumps were performed at 064/1658Z. The recovery load was patched to allow MTS failure flags to remain at 064/1704Z. At 064/1707Z, telemetry diagnostics were performed to monitor MTS current analogs. Ran a fine attitude determination for SWSP at 064/1753Z. Performed a attitude reference update at 064/1857Z. Enabled TRSWCC to recover from SWSP at 064/1929Z. Executed a real-time slew to the intercept attitude at 064/1931Z. Performed another fine attitude determination at 064/2041Z. Recovery from safemode was completed at 064/2100Z. The Health and Safety (H&S) Science Mission Specification loads began execution at 064/210000Z. These loads will continue until 065/220000Z when normal science commanding will resume. After transition to the H&S loads, the following ROPs were performed: SR-1A - twice clearing the SSR EDAC error counter. DF-18A - at 064/2115Z to dump and reset the engineering status buffer. IC-02 - SSA transmitter #1 was turned on for the interval 064/2318Z - 064/2344Z. NS-03 - at 064/2350Z to dump and reset the NSSC-1 status buffer. DF-07A - the ephemeris load was uplinked at 065/0104Z. IC-02 - SSA transmitter #2 was turned on at 065/0104Z and turned off at 065/0123Z 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /CAW